OpenStudy (anonymous):

find the axis of symmetry for this parabola y=-5x^2-10x-13

OpenStudy (anonymous):

write the answer as an equation

OpenStudy (irishboy123):

complete the square \(y=-5x^2-10x-13\) might be easier to do it starting with this \(y=-5[x^2+2x+\frac{13}{5}]\)

OpenStudy (anonymous):

-1?

OpenStudy (irishboy123):

\(\large \checkmark\)

OpenStudy (anonymous):

It's not that

OpenStudy (irishboy123):

it is x = -1 if you copied the question correctly

OpenStudy (anonymous):

I copied the question right but -1 wasn't the answer

OpenStudy (anonymous):

try this one y=-x^2+6x-8

OpenStudy (loser66):

@Kaelyn78 you want us do your homework, don't you?

OpenStudy (irishboy123):

you could do that one in your head!

OpenStudy (anonymous):

these are practice problems, not homework

OpenStudy (anonymous):

this paper is not graded

OpenStudy (anonymous):

and no I can't, I don't know how to find the axis of symmetry in my head

OpenStudy (irishboy123):

we still haven't finished the first one if the answer is not x = -1, what is it?!

OpenStudy (anonymous):

I don't know, that's why I posted it here.

OpenStudy (anonymous):

1 or -1

OpenStudy (loser66):

The first one, @IrishBoy123 showed you all the steps + proof. You said it is not the right answer without any reason. I am with him and don't get why you go to other one without finish the first one. :)

OpenStudy (irishboy123):

do it another way \(y=-5x^2-10x-13\) \(y' = -10 x - 10\) \(y' = 0 \implies x = -1\)

OpenStudy (anonymous):

because I entered that answer into the computer and it said it isn't right

OpenStudy (astrophysics):

Irish boy is right!

OpenStudy (anonymous):

Computers don't lie, and they don't give you an explanation of why it's wrong

OpenStudy (anonymous):

I came to the same conclusion he had, I agree with him, but the computer doesn't.

OpenStudy (astrophysics):

You can check another way by finding the vertex, \[x= -\frac{ b }{ 2a }\]

OpenStudy (astrophysics):

\[x = - \frac{ (-10) }{ 2(-5) } = -1\]

OpenStudy (anonymous):

Guys they don't give you an option to do the problem a second time. I entered the answer and it was wrong so I just have to move on to the next question. Proving it to me over and over again will not change anything.

OpenStudy (anonymous):

If you ever meet the developers of the program I'm working with you can argue with them about that. Thank you all for your help on that problem.

OpenStudy (irishboy123):

put each new question in a new thread and let's see if we can sort out that pesky computer of yours. perhaps you should wire it to a lie detector :p

OpenStudy (astrophysics):

lool

OpenStudy (anonymous):

That would be great... Technology is great when it works, horrible when it doesn't